OpenStudy (anonymous):

u = 6i + 7j − k. Find a unit vector in the direction of u

OpenStudy (anonymous):

\[unit ~vector~u=\frac{ vector~u }{\left| vector~u \right| }\] \[\left| u \right|=\sqrt{6^2+7^2+(-1)^2}\]

OpenStudy (anonymous):

i need answer in vector form < 1, 2 ,3 >

OpenStudy (accessdenied):

The magnitude of the vector is just a scalar, so dividing by the magnitude is the equivalent of dividing each component by the magnitude.

OpenStudy (anonymous):

so divide 6, 7, and -1 by the magnitude

OpenStudy (kainui):

@Parth1994 Yeah. Your new vector will have a length of 1 if you want to check it yourself to make sure. After all, it should be a unit vector right?

OpenStudy (anonymous):

thanks got it!! appreciate it

OpenStudy (compassionate):

@April115